OpenStudy (anonymous):

f(x) = 4throot(x), c = 1; find the slope and the y-intercept of the tangent line T.

OpenStudy (turingtest):

could you tell me where you're having trouble?

OpenStudy (turingtest):

first you need the derivative, so let me know what you get for that

OpenStudy (anonymous):

f(x) = 4throot(x) = x^(1/4); f'(x) = (1/4)x^(-3/4) = (1/4)(1/[x^(3/4)]) = 1/(4x^[3/4]).....

OpenStudy (turingtest):

and the value at x=1 (you really shouldn't say c=1, that is just the usage in your book, it may confuse people)

OpenStudy (turingtest):

*I mean and what is the value at x=1?

OpenStudy (anonymous):

right it's 1

OpenStudy (turingtest):

f'(1)=1 ??? I don't think so, check again...

OpenStudy (anonymous):

i mean 4!!!!

OpenStudy (anonymous):

I MEAN 1/4!!!!!!!

OpenStudy (anonymous):

...